OpenStudy (anonymous):

Sin(arctan(-0.2)) Explanation?

OpenStudy (debbieg):

Do you know what the functions sine and arctangent (inverse tangent) mean, and where to find them on your calculator? This is not one of the "common" unit circle angle values, so you would need to use a calculator to evaluate. You could either enter it all in one step just as it is above; or first evaluate \(\large \tan^{-1}(-0.2)\) and then evaluate the sine of the result.

OpenStudy (anonymous):

Is there a way you could do it without a calculator? Using a unit or circle or something?

OpenStudy (debbieg):

Do you think you're supposed to it without a calculator? hmmm.........

OpenStudy (anonymous):

Unfortunately my teacher thinks I am supposed to

OpenStudy (debbieg):

Well, lessee.... maybe I'm missing something. We need, first, \(\large \tan^{-1}(-0.2)=\tan^{-1}(-\dfrac{2}{10})=\tan^{-1}(-\dfrac{1}{5})\) So that is the Q4 angle x such that: \(\Large \tan(x)=-\dfrac{1}{5}\) So the point (1,-5) is on the terminal side. a HAH!! Now I see it......

OpenStudy (debbieg):

do you see how you can find the sine of this angle? :) 

OpenStudy (anonymous):

Haha. Uhhhhh...

OpenStudy (debbieg):

It's actually a clever little problem. I didn't see it until I started to "think out loud" about what we know, from the given information. The beauty of it is, we don't NEED to KNOW what the ANGLE is to find it's sine value. We can use the definition of sine as it relates to a point on the terminal side of the angle. And we have that, since we know the tangent of the angle, that we want the sine of! :)

OpenStudy (anonymous):

I'm really trying to understand...

OpenStudy (debbieg):

Do you know the equations for the trig functions, using a point on the terminal side of the angle? \(\large \sin\alpha=y/r\) where \(\large r=\sqrt{x^2+y^2}\) Does that look familiar?

OpenStudy (anonymous):

Yes

OpenStudy (debbieg):

Actually, maybe I should back up. Do you understand the picture I drew above? how I got the point on the terminal side of the angle, using the fact that: We want: \(\large \sin x\) where: \(\large x=\tan^{-1}(-0.2)\) Do you follow how I got the diagram of the angle that I drew, from that?

OpenStudy (anonymous):

Yes because Tangent is y/x so you just plugged in 1 and 5, right?

OpenStudy (anonymous):

-5*

OpenStudy (debbieg):

Exactly. Once the decimal was converted to a reduced fraction, you can use that ratio to give you a point on the terminal side.

OpenStudy (anonymous):

And how does that lead us to Sine?

OpenStudy (debbieg):

(actually, even unreduced fraction would have worked - I would just have used the point (2, -10) in that case.) OK, here is how that gets us to sine: we want the sine of THAT angle, the angle that I drew above. We know that \(\large \sin\alpha=y/r\) where \(\large r=\sqrt{x^2+y^2}\) So now, just compute r using that point (1, -5), and then use y and r to find the sine of the angle! :)

OpenStudy (debbieg):

Once you have a point on the terminal side of an angle, that's all you need - you can get all 6 of the trig functions from that point.

OpenStudy (anonymous):

So I find r using the pythagorean theorem, plug that into "-5/r" and find the sine of that fraction?

OpenStudy (debbieg):

When you plug r into -5/r, that IS the SINE of that ANGLE. There is nothing else to do.

OpenStudy (anonymous):

Can you solve it and see if we get the same answer?

OpenStudy (debbieg):

\(\large \sin\alpha=y/r\) where \(\large r=\sqrt{x^2+y^2}\) and (x,y) is a point on the terminal side of \(\alpha\). You have the point, (x,y). All that's left is to compute r and then plug it into \(\large \sin\alpha=y/r\). Yes, I've already solved it. Let me know what you get, I'll tell you if you are right. :)

OpenStudy (anonymous):

I got \[(-5\sqrt{26})/26\] Is that right?

OpenStudy (debbieg):

Yes, that's it! And you even rationalized the denominator - good for you. :)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (debbieg):

You're welcome! and welcome to Open Study! :)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Hopefully you still see this, why could the point on the unit circle not be (-1,5) instead of (1,-5)?

OpenStudy (debbieg):

Ah, great question! Here is why: You might recall, when you learned about the inverse functions, you (hopefully, lol) learned that each one has a SPECIFIC and defined range to go with it. E.g., the function y=tan(x) is, in general NOT an invertible function because it isn't one-to-one (you can get the same tangent value for multiple values of the input, x). How do we get around that, and make it so that we can have inverse functions? Why, we restrict the domain of the function y=tan(x) over which we will invert it! That way, the inverse function IS, well, a function (because to be a function, it must be true that every value in the domain can map to ONLY ONE value in the range. So by restricting the range of the original function that we are inverting, we restrict the range of the inverse. So for y=arctan(x), the domain of y=tan(x) is restricted to \(-\pi/2<x<\pi/2\), in other words, quadrants 4 and 1. Thus, the range of y=arctan(x) is \(-\pi/2<x<\pi/2\). So, y=arctan(x) MEANS "that angle y in Q4 or Q1 that has tan(y)=x". Since the value here was negative, we know that we are in Q4.

OpenStudy (anonymous):

Ahhhhhh, right. So we use the domain of the inside trig function? So in this problem it would be tangent instead of sine?

OpenStudy (anonymous):

So if I had:\[\tan (\cos^{-1} (-12/13)\] it would be in the second quadrant with a point at (-12,13)?

OpenStudy (debbieg):

Yes, that's exactly right, because the range of the inverse cosine function is 0 to pi (Q1 and Q2), so arccos(-12/13) is an angle in Q2. :)

OpenStudy (anonymous):

Sorry I meant a point at (-12,5) with a radius of 13

OpenStudy (anonymous):

Or was I right the first time?